OpenStudy (erinweeks):

Write the equation in slope-intercept form. What are the slope and y-intercept? -11x + 9y = -12

OpenStudy (erinweeks):

A. y = 11/9x - 4/3 slope: 11/9 y-intercept: 4/3 B. y = -11/9x + 4/3 slope: 11/9 y-intercept: 4/3 C. y = 11/9x - 4/3 slope: 11/9 y-intercept: -4/3 D. y = 11/9x + 4/3 slope: 4/3 y-intercept: 11/9

OpenStudy (anonymous):

slope intercept form is y=mx +c here

OpenStudy (anonymous):

here we have -11x +9y=-12

OpenStudy (anonymous):

or 9y=11-12 or y=11/9 x -12/9 y=11/9 x -4/3

OpenStudy (anonymous):



OpenStudy (anonymous):

here slope =11/9 and y-intercept is -4/3

OpenStudy (erinweeks):

thank you so much